\vspace{6mm}

 \begin{thebibliography}{55}

\vspace{4mm}

\bibitem{Beckner}
W. Beckner, Inequalities in Fourier analysis, \emph{Ann. of Math.},
102 (1975), 159--182.


\bibitem{Bennett Sharpley1988}
C. Bennett and R. Sharpley, {\it Interpolation of Operators}.
Academic Press, New York, 1988.

\bibitem{Brascamp}
H.J. Brascamp and  E.H. Lieb, Best constants in Young's inequality,
its converse, and its generalization to more than three functions,
\emph{Advances in Math.}, 20(2) (1976), 151--173.

\bibitem{Capone-Formica-Giova-2013}
C. Capone, M.R. Formica and R. Giova, Grand Lebesgue spaces with
respect to measurable functions, \emph{Nonlinear Anal.}, 85 (2013),
125--131.

\bibitem{Capone-Fiorenza-2005-small}
C. Capone and A. Fiorenza, On small Lebesgue spaces, \emph{J. Funct.
Spaces Appl.} 3 (2005), no. 1, 73--89.

 \bibitem{Cigler}
J. Cigler, Normed ideals in $L^1(G)$,\emph{ Indag. Math.}, 31
(1969), 273--282.

\bibitem{Ermakov}
V. Ermakov and E.I. Ostrovsky, \emph{Continuity Conditions,
Exponential Estimates, and the Central Limit Theorem for Random
Fields}, Moscow, VINITY, (1986), (in Russian).


\bibitem{Fiorenza-2000-duality}
 A. Fiorenza, Duality and reflexivity in grand Lebesgue spaces.
Collect. Math. 51 (2) (2000), 131--148.


 \bibitem{Fiorenza-Formica-Nonlinear_Anal_2018}
A. Fiorenza, M.R. Formica, A. Gogatishvili, T. Kopaliani and J.M.
Rakotoson, Characterization of interpolation between grand, small or
classical Lebesgue spaces, \emph{Nonlinear Anal.}, 177 (2018),
422--453.

\bibitem{fioformicarakodie2017}
A. Fiorenza, M.R. Formica and J.M. Rakotoson, Pointwise estimates
for {$G\Gamma$}-functions and applications, \emph{Differential
Integral Equations} 2017; 30 (11-12): 809--824.

\bibitem{Fiorenza-Karadzhov-2004}
A. Fiorenza and G.E. Karadzhov, Grand and small Lebesgue spaces and
their analogs, \emph{Z. Anal. Anwendungen}, 23 (4) (2004), 657--681.

\bibitem{Fiorenza-Mercaldo-Rakotoson-2001}
A. Fiorenza, A. Mercaldo and J.M. Rakotoson, Regularity and
comparison results in grand Sobolev spaces for parabolic equations
with measure data. \emph{Appl. Math. Lett.}, 14 (2001), no. 8,
979--981.

\bibitem{Fiorenza-Mercaldo-Rakotoson-2002}
A. Fiorenza, A. Mercaldo and J.M. Rakotoson, Regularity and
uniqueness results in grand Sobolev spaces for parabolic equations
with measure data. \emph{Discrete Contin. Dyn. Syst.}, 8 (2002), no.
4, 893--906.

\bibitem{Formica-Giova-2015-Boyd}
M.R. Formica and R. Giova, Boyd indices in generalized grand
Lebesgue spaces and applications, \emph{Mediterr J. Math.}, 12 (3)
(2015), 987--995.

\bibitem{Fournier}
J.F. Fournier, Sharpeness in Young's inequality for convolution,
\emph{Pacific J. Math.}, 72 (1977), no. 2, 383--397.

\bibitem{Greco-Iwaniec-Sbordone-1997}
L. Greco, T. Iwaniec and C. Sbordone, Inverting the $p$-harmonic
operator, \emph{Manuscripta Math.}, 92 (1997), no. 2, 249--258.

\bibitem{Gurkanli1}
A.T. G\"{u}rkanli, Inclusions and the approximate identities of the
generalized grand Lebesgue spaces, \emph{Turk. J. Math.}, 42 (6)
(2018), 3195--3203.

\bibitem{Gurkanli2}
A.T. G\"{u}rkanli, Multipliers of some Banach ideals and
Wiener-Ditkin sets, \emph{Math. Slovaca} 55 (2005), no. 2, 237--248.


\bibitem{Turan}
A.T. G\"{u}rkanli, Multipliers of grand and small Lebesgue Spaces,
arXiv:1903.06743v1 [math.FA] 15 Mar 2019.


\bibitem{Hardy}
G.H. Hardy, J.E. Littlewood and G. Polya, \emph{Inequalities},
Cambridge University Press, 1952.

\bibitem{Hewitt1970}
E. Hewitt and K. Ross, \emph{Abstract Harmonic Analysis Vol. II:
Structure and analysis for compact groups. Analysis on locally
compact Abelian groups}, Springer-Verlag, New York-Berlin, 1970.

\bibitem{Hewitt}
E. Hewitt and K. Ross, \emph{Abstract Harmonic Analysis vol. I:
Structure of Topological Groups. Integration Theory, Group
Representations}, Springer-Verlag, Berlin, 1969.

\bibitem{Iwaniec-Sbordone-1992}
T. Iwaniec and C. Sbordone, On the integrability of the Jacobian
under minimal hypotheses. \emph{Arch. Rat.Mech. Anal.}, 119, (1992),
129--143.

\bibitem{Iwaniec-Koskela_Onninen-2001}
T. Iwaniec, P. Koskela and J. Onninen, Mapping of finite distortion:
monotonicity and continuity. \emph{Invent. Math.}, 144 (2001),
507--531.


\bibitem{Jain-Singh-Singh_Stepanov-2019}
P. Jain, A.P. Singh, M. Singh and V.D. Stepanov, Sawyer's duality
principle for grand Lebesgue spaces, \emph{Math. Nachr.}, 292(4)
(2019), 841--849.

\bibitem{Jain-Jain-2017}
P. Jain, S. Jain, Generalized convolution inequalities and
application. \emph{Mediterr. J. Math.}, 14 (2017),  no. 4, Art. 159,
13 pp.

\bibitem{Jain-Jain-Kumar-2015}
P. Jain, S. Jain and K. Kumar, On fractional convolutions and
distributions. \emph{Integral Transforms Spec. Funct.}, 26 (2015),
no. 11, 885--899.



\bibitem{KozOs}
Yu V. Kozachenko and E.I. Ostrovsky, The Banach Spaces of random
variables of sub-Gaussian type (in Russian). \emph{Probab. and Math.
Stat.}, Kiev, KSU, 32 (1985), 43--57.


\bibitem{KozOsSir2017}
Yu V. Kozachenko, E.I. Ostrovsky and L. Sirota, Relations between
exponential tails, moments and moment generating functions for
random variables and vectors. arXiv:1701.01901v1 [math.FA], 8 Jan
2017.


\bibitem{LiflOstSir}
E. Liflyand, E. Ostrovsky and L. Sirota, Structural properties of
bilateral grand Lebesque spaces, \emph{Turkish J. Math.} 34 (2)
(2010), 207--219.

\bibitem{Molchanova2019}
A. Molchanova, A note on the continuity of minors in grand Lebesgue
spaces, \emph{J. Fixed Point Theory Appl.}, 21(2),  Article number
49, (2019), DOI: 10.1007/s11784-019-0686-y.

\bibitem{Naimark}
M.A. Naimark, Normed rings, \emph{Bull. Amer. Math. Soc.}, 69 (2)
(1963), 193--195.


\bibitem{Ostrovsky1}
E. Ostrovsky, \emph{Exponential estimates for random fields and its
applications}, Moscow - Obninsk: OINPE, 1999 (Russian).

\bibitem{Ostrovsky2}
E.I. Ostrovsky and L. Sirota, Moment Banach spaces: theory and
applications, HIAT Journal of Science and Engineering, C, vol.4,
issues 1-2, (2007), 233--262.

\bibitem{Ostrovsky3}
E.I. Ostrovsky and L. Sirota, Multidimensional Dilation Operators,
Boyd and Shimogaki indices of Bilateral Weight Grand Lebesgue
Spaces. arXiv:0809.3011[math.FA] 17 Sep 2008.

\bibitem{Ostrovsky4}
E.I. Ostrovsky and L. Sirota, Fundamental functions for Grand
Lebesgue Spaces, arXiv:1509.03644v1 [math.FA] 11 Sep 2015.

\bibitem{Ostrovsky5}
E.I. Ostrovsky and L. Sirota, Boundedness of operators in bilateral
Grand Lebesgue Spaces, with exact and weakly exact constant
calculation. arXiv:1104.2963v1 [math.FA] 15 Apr 2011.

\bibitem{Talenti}
G. Talenti, Inequalities in Rearrangement Invariant Function Spaces,
\emph{Nonlinear Analysis, Function Spaces and Applications}.
\emph{Prometheus}, Prague, 5, (1995), 177--230.

\bibitem{Zelazko}
W. Zelazko, On the algebras $L_p$ of locally compact groups. \emph{
Colloq. Math.}, 8 (1) (1961), 115--120.


\vspace{5mm}


 \end{thebibliography}

\documentclass[10pt]{article}
\usepackage[b5paper]{geometry}
\usepackage{amsmath,amsthm,amsfonts,amssymb,MnSymbol,mathrsfs,latexsym}
\usepackage[english]{babel}

\usepackage[applemac]{inputenc}
\usepackage[all]{xy}
\numberwithin{equation}{section}

\DeclareMathOperator*{\esssup}{ess\,sup}

  \newcommand{\const}{\rm const}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\theoremstyle{theorem}

\newtheorem{remark}{Remark}[section]

\renewenvironment{proof}{{\bf{Proof.}}}{\hfill $\Box$ \\}

\begin{document}

\title{ \large \textbf{Grand Lebesgue Spaces are really Banach algebras relative to the convolution on unimodular locally compact groups}}

\footnotesize\date{}

\author{\normalsize \textbf{Maria Rosaria FORMICA ${}^{1}$},   \normalsize\textbf{Eugeny OSTROVSKY
${}^2$} and \normalsize\textbf{Leonid SIROTA ${}^3$}}

\maketitle

\begin{center}
{\footnotesize ${}^{1}$ Universit\`{a} degli Studi di Napoli \lq\lq Parthenope\rq\rq, via Generale Parisi 13,\\
Palazzo Pacanowsky, 80132,
Napoli, Italy.} \\

\vspace{2mm}

{\footnotesize e-mail: mara.formica@uniparthenope.it} \\

\vspace{4mm}

{\footnotesize ${}^{2,\, 3}$  Bar-Ilan University, Department of Mathematics and Statistics, \\
52900, Ramat Gan, Israel.} \\

\vspace{2mm}

{\footnotesize e-mail: eugostrovsky@list.ru}\\

\vspace{2mm}

{\footnotesize e-mail: sirota3@bezeqint.net} \\

\vspace{4mm}

\end{center}

\begin{abstract}
We prove that the Grand Lebesgue Space, builded on a unimodular
locally compact topological group, forms a Banach algebra relative
to the convolution.
\end{abstract}


\vspace{4mm}

{\bf Keywords:} Grand Lebesgue Spaces, convolution, Haar measure,
Young inequality, Lebesgue-Riesz space, locally compact topological
groups, unimodular group, modulus of continuity, Beta-function.

\vspace{2mm} \noindent {\it 2010 Mathematics Subject
Classification}:
46E30,   
44A35,     
43A10.    

\vspace{5mm}

\section{Definitions. Statement of problem. Notations.}

\vspace{5mm}

Let $G$ be a unimodular locally compact topological group equipped
with bi-invariant Borelian Haar measure $\mu$. Define, as usually,
the convolution between two measurable integrable functions $ \
f,g:G \to R \ $ by
$$
(f*g)(x) \stackrel{def}{=} \int_G f(y) g(y^{-1} \, x) \, d\mu(y).
$$

The usual Lebesgue-Riesz spaces $L_p = L_p(G)=L_p(G,\mu)$ are the
spaces of all measurable functions $f: G\to \mathbb R$ defined by
the norms
$$
 \begin{array}{ll}
      |f|_p := \displaystyle \left[ \int_G |f(x)|^p
\, d\mu(x) \right]^{1/p}, & 1\leq p<\infty \\ \\
|f|_{\infty} := \displaystyle \esssup_{x \in G} |f(x)|, & p=\infty
    \end{array}
$$

The well-known Young (or Hausdorff-Young) inequality has the form

\begin{equation} \label{HausYoung}
|f*g|_r \le |f|_p \ |g|_q, \ 1 \le p,q,r \le \infty,
\end{equation}
where
\begin{equation} \label{restiction}
 1+1/r = 1/p + 1/q,
\end{equation}
see e.g. \cite{Hardy}, \cite{Zelazko}, \cite{Bennett Sharpley1988}
and \cite{Beckner}, \cite{Brascamp}, \cite{Fournier} for the
computation of the best possible constant in (\ref{HausYoung}).\par

\noindent Generalized convolution inequalities and applications have
been obtained in \cite{Jain-Jain-2017} and fractional convolutions
in \cite{Jain-Jain-Kumar-2015}.

Choosing $r = p, \, q = 1$ in (\ref{HausYoung}), we deduce
\begin{equation} \label{HausYoungcomp}
|f*g|_p \le |f|_p \ |g|_1.
\end{equation}

If in addition the group $G$ is also compact, one can suppose
$\mu(G)
= 1$ so that 
$|g|_1 \le |g|_p$. Therefore
\begin{equation} \label{HausYoungBanAlg}
|f*g|_p \le |f|_p \ |g|_p, \  p \in [1,\infty].
\end{equation}

The last inequality shows that the space $L_p(G,\mu)$, relative to
the convolution bilinear operator, forms in this case a {\it Banach
algebra}, see \cite{Zelazko}.

\vspace{2mm}

Our aim is to extend inequality \eqref{HausYoungBanAlg} to the
so-called Grand Lebesgue Spaces, improving some results recently
obtained in \cite{Turan}; see also \cite{Gurkanli1},
\cite{Gurkanli2}. \par

 \vspace{2mm}

In more details, we will find a sufficient condition on the
generating function of these spaces under which they form also a
 {\it Banach algebra}, relative to the convolution. \par

\begin{center}

 \vspace{2mm}
\ {\sc Brief note about  Grand Lebesgue Spaces. } \\

\end{center}

\vspace{2mm}

 We recall here, for reader convenience, some known definitions and facts about the theory of the Grand Lebesgue Spaces which we will use.
 We consider only the case of the space on the set $G$ equipped  with the {\it Haar} measure $\mu$. \par

 Let $\psi = \psi(p) = \psi[b](p), \ p \in [1,b), \ b = \const \in (1,\infty]$, be a numerical valued non-negative function, not necessarily finite in every point, such that

\begin{equation} \label{Positive psi}
 \inf_{p \in [1,b)} \psi(p) > 0
\end{equation}
and

\begin{equation} \label{Positive norming}
  \psi(1) \in( 0, \ \infty).
\end{equation}
The set of all such functions will be denoted by $ \Psi = \Psi[b] =
\{\psi(\cdot) \}$. If $b < \infty$, one can assume $p \in [1,b]$.
\par

\vspace{2mm}

We will assume, without loss of generality, \, $\psi(1) = 1$ for all
such functions $\psi(\cdot)$.

 \vspace{2mm}

\noindent For instance, the following functions
$$
\psi_{(m)}(p) := p^{1/m}, \ \ m > 0, \  \ p \in [1,\infty)
$$
or
$$
\psi^{(b; \beta)}(p) := (b-1)^{\beta} (b-p)^{-\beta}, \ \ p \in
[1,b), \  \ \beta \ge 0, \ \ b\in (1,\infty)
$$
satisfy \eqref{Positive psi} and \eqref{Positive norming}, with
$\psi(1) = 1$.

\vspace{2mm}

The (Banach) Grand Lebesgue Space $G\psi =G\psi [b]$ consists of all
the real (or complex) numerical valued measurable functions $f: G
\to R$ having finite norm, defined by

 \begin{equation} \label{norm psi}
    || f ||_{G\psi} = ||f||_{G\psi[b]} \stackrel{def}{=} \sup_{p \in [1,b)} \left[ \frac{|f|_p}{\psi(p)} \right].
 \end{equation}

 \vspace{4mm}

 The function $\psi = \psi(p)=\psi[b](p) $ is named as the {\it generating function} for this space. \par

If for instance
$$
 \psi(p) = \psi^{(r)}(p)=\left\{
  \begin{array}{ll}
    1  & , \ \ p=r \\
    +\infty &  , \ \ p\neq r
  \end{array}
\right.
$$
where $ r\in [1,\infty),  \ C/\infty := 0, \ C \in \mathbb R$,
(extremal case), then the corresponding $ G\psi^{(r)}(p)$ space
coincides with the classical Lebesgue-Riesz space $L_r = L_r(G)$.

\vspace{4mm}

Note that the space $L^{b),\theta}$ introduced in
\cite{Greco-Iwaniec-Sbordone-1997} (see \cite{Iwaniec-Sbordone-1992}
for $\theta =1$), through the norm

$$
||f||_{L^{b),\theta}}=||f||_{b, \theta,G} \stackrel{def}{=} \sup_{0
< \epsilon \le b - 1} \left[ \ \epsilon^{\theta/(b-\epsilon)} |f|_{b
- \epsilon} \ \right], \ \theta \ge 0
$$
quite coincides, up to equivalence of the norms, with the
appropriate one $G\psi^{(b; \beta)}$, with $\beta = \theta/b$.

Namely, 
$$
C^-(b,\theta,G)||f||_{G\psi^{(b; \beta)}}\leq ||f||_{L^{b),\theta}}
\leq C^+(b,\theta,G)||f||_{G\psi^{(b; \beta)}},
$$
where $0<C^-(b,\theta,G)\leq C^+(b,\theta,G)<\infty$.
 \vspace{4mm}

Let $f: G \to R$ be a measurable function and $b \in (1,\infty]$
such that
$$
|f|_p < \infty,  \ \ \forall p \in [1,b);
$$
the so-called {\it natural function} $\psi^{(f)}(p)$ for $f$ is
defined by
$$
\psi^{(f)}(p) \stackrel{def}{=} |f|_p.
$$
Obviously,
$$
||f||_{G\psi^{(f)}} = 1.
$$

\vspace{4mm}

 The Grand Lebesgue Spaces have been widely investigated, e.g.
\cite{Fiorenza-2000-duality}, \cite{Capone-Fiorenza-2005-small},
\cite{Capone-Formica-Giova-2013}, \cite{Formica-Giova-2015-Boyd},
\cite{Iwaniec-Koskela_Onninen-2001}, \cite{KozOs},
\cite{LiflOstSir}, \cite{Ostrovsky1} - \cite{Ostrovsky5},
\cite{Jain-Singh-Singh_Stepanov-2019}, \cite{Molchanova2019}, etc.
They play an important role in the theory of Partial Differential
Equations (PDEs) (see e.g. \cite{Fiorenza-Mercaldo-Rakotoson-2001},
\cite{Fiorenza-Mercaldo-Rakotoson-2002},
\cite{Greco-Iwaniec-Sbordone-1997}, \cite{fioformicarakodie2017}),
in interpolation theory (see e.g. \cite{Fiorenza-Karadzhov-2004},
\cite{Fiorenza-Formica-Nonlinear_Anal_2018}), in the theory of
Probability (\cite{Ermakov}, \cite{KozOsSir2017},
\cite{Ostrovsky3}-\cite{Ostrovsky5}), in Statistics and in the
theory of random fields (see e.g. \cite{KozOs}, \cite[chapter
5]{Ostrovsky1}), in Functional Analysis and so on.
\par

 These spaces are rearrangement invariant (r.i.) Banach functional spaces; its fundamental function has been studied in  \cite{Ostrovsky4}. They
 not coincides, in the general case, with the classical rearrangement invariant spaces: Orlicz, Lorentz, Marcinkiewicz, etc., see \cite{LiflOstSir}, \cite{Ostrovsky2}.

\vspace{5mm}

\section{Main result.}

\begin{theorem}\label{main result}

Let $G$ be a unimodular locally compact topological group equipped
with bi-invariant Haar measure $\mu$ and let $G\psi $ be the
(Banach) Grand Lebesgue space builded on $G$. If $ \psi(1) = 1$,
then $G\psi$ is also a Banach algebra under the convolution as a
multiplicative operation, that is
\begin{equation} \label{result}
||f * g||_{G\psi} \le ||f||_{G\psi} \cdot ||g||_{G\psi}
\end{equation}
for all $f,g: G\to \mathbb R$ in $G_{\psi}$.
\end{theorem}

\vspace{4mm}

\begin{proof}
 Let $f,g$ be two non-zero functions from the space $ G\psi$;  we assume, without loss of generality, that
$$
||f||_{G\psi} = ||g||_{G\psi} = 1.
$$
From the definition of the norm in the Grand Lebesgue Spaces it
follows immediately that
$$
|f|_p \le \psi(p), \ \ |g|_p \le \psi(p), \  \ p \in [1,b).
$$
We apply  once again the Young inequality (\ref{HausYoungcomp})
$$
|f*g|_p \le |f|_p \, |g|_1 \le \psi(p) \, |g|_1,
$$
thus
$$
\frac{|f*g|_p}{\psi(p)} \le |g|_1 \le \psi(1) = 1
$$
and, taking the supremum over $p \in [1,b)$, we have

$$
\sup_{p \in [1,b)} \left\{ \ \frac{|f*g|_p}{\psi(p)} \ \right\}
\le |g|_1 \le \psi(1) = 1.
$$

Therefore

\begin{equation*} 
||f*g||_{G\psi} \le 1 = ||f||_{G\psi} \cdot ||g||_{G\psi}.
\end{equation*}

\end{proof}

\vspace{2mm}

\begin{remark}
 In the next section we will see that equality in \eqref{result} can be achieved.
\end{remark}

\section{An example.}

\vspace{5mm}

 Let now  $ G =\mathbb R$ with the classical Lebesgue measure $d \mu = dx$ and the ordinary convolution
\begin{equation}\label{convolution in R}
(f*g)(x)=\int_{-\infty}^{+\infty} f(y)g(x-y)\, dy,
\end{equation}
for all measurable integrable functions $f,g$ defined on $\mathbb
R$.

Define the Gaussian  density
\begin{equation}\label{gaussian_density}
z =z(x) = z_{\sigma}(x) := (\sigma \ \sqrt{2 \ \pi})^{-1} \, \exp
\left\{ \frac{-x^2}{2\sigma^2}  \right\}, \ \ \sigma = \const > 0;
\end{equation}
then its natural function $\psi^{(z,\sigma)}(p)$ has the form

\begin{equation}\label{gaussian_psi}
\psi_{\sigma}(p)= \psi^{(z_{\sigma})}(p) = | \ z_{\sigma}(\cdot) \
|_p = (2 \pi)^{ 1/(2p) - 1/2 } \ p^{-1/(2p)} \sigma^{1/p - 1}, \ \ p
\in [1,\infty).
\end{equation}

One can define formally the value $\psi^{(z_{\sigma})}(p)$ for $p
= \infty$ as

$$
\psi_{\sigma}(\infty)  \stackrel{def}{=} \lim_{p \to \infty} \psi_{\sigma}(p) = (\sigma \ \sqrt{2 \ \pi})^{-1} = \max_{x \in R} z_{\sigma}(x).
$$

 Notice that $\psi_{\sigma}(1) = 1$. By virtue of Theorem \ref{main result} we conclude that the Grand Lebesgue Space $G\psi^{(z_{\sigma})}(p)$, builded over
 the ordinary real line $ \mathbb R$, relative to the convolution operation, forms also a Banach algebra, despite the (commutative) group $(\mathbb R,+)$ is not compact.

\vspace{2mm}

Moreover, let us choose $f(x) = g(x) = z_{1}(x)$, then

$$
||f||_{G\psi_{1}} = ||g||_{G\psi_{1}} = 1.
$$
Furthermore,
$$
h(x) := (f*g)(x)= z_{\sqrt{2}}(x);
$$

$$
|h|_p = (2 \pi)^{ 1/(2p) - 1/2 } \ p^{-1/(2p)} [\sqrt{2}]^{1/p - 1}, \ p \in [1,\infty),
$$
so that

$$
  \frac{|h|_p}{\psi_{1}(p)} = [\sqrt{2}]^{1/p - 1},
$$
and

$$
||h||_{G\psi_{1}}  = \sup_{p \ge 1} [\sqrt{2}]^{1/p - 1} = 1.
$$
 Thus, in this case, we have

$$
||h||_{G\psi_{1}} = ||f*g||_{G\psi_{1}}  = ||f||_{G\psi_{1}} \cdot
||g||_{G\psi_{1}}.
$$

\vspace{2mm}

 Note that the case $ \ G = \mathbb R^d \ $ may be investigated analogously.

\vspace{5mm}

\section{Concluding remarks.}

\vspace{5mm}

 \ {\bf A.} Let $E$ be an arbitrary translation invariant Banach functional space  with norm $||\cdot||_E $ builded over $ \ G \ $
 and such that

$$
\forall g \in L_1 \ \Rightarrow |g|_1 \le ||g||_E.
$$
 We have, using the triangle inequality,

\begin{equation*}
\begin{split}
||f*g||_E = & || \int_G f(y^{-1} x) \ g(y) \ \mu(dy)  ||_E \le
\int_G
||f||_E \ |g(y)| \ \mu(dy) \\
= &||f||_E \ |g|_1 \le    ||f||_E \cdot ||g||_E.
 \end{split}
\end{equation*}
 Therefore, $E$ is again a Banach algebra.

\vspace{4mm}

{\bf B.} Let us prove the {\it necessity} of the condition
\eqref{restiction} at least in the case $G= \mathbb R^d, \ d =
1,2,3,\ldots$. Suppose that for {\it arbitrary} non-zero and
non-negative numerical valued smooth functions $f,g$, defined on
$\mathbb R^d$ and having compact support, it is

\begin{equation} \label{estimat}
|f*g|_r \le |f|_p \ |g|_q, \ \   p,q,r \ge 1.
\end{equation}
Then $ \ 1+ 1/r = 1/p + 1/q$. \par

Actually, one can use the so-called scaling method, see e.g.
\cite{Talenti}. Assume that (\ref{estimat}) is satisfied. Define for
any positive constant $ \lambda > 0 $ the dilation operator

$$
T_{\lambda}[f](x) = f(\lambda \ x),  \ \ x \in \mathbb R^d.
$$
We deduce from \eqref{estimat}

\begin{equation} \label{estim}
|T_{\lambda}f*T_{\lambda}g|_r \le |T_{\lambda}f|_p \
|T_{\lambda}g|_q, \  \  p,q,r \ge 1
\end{equation}
and, after simple calculations,

$$
\lambda^{-d - d/r} \ |f*g|_r \le \lambda^{-d/p - d/q} \ |f|_p \ |g|_q.
$$
Since $\lambda$ is a arbitrary positive value, we conclude $ \  -d -
d/r = -d/p - d/q,  \ $ which is quite equally to
\eqref{restiction}.\par

\vspace{5mm}

 {\bf C.} As long as the Grand Lebesgue Space  $ \ G\psi_{1}$ forms a Banach algebra, it is interesting,
by our opinion, to consider its ideals, maximal or not. We apply
here well-known classical results from \cite{Cigler}, \cite{Hewitt},
\cite{Hewitt1970}, \cite{Naimark}.\par

 Define as usually for this purpose the Fourier transform for every function from this space

\begin{equation} \label{Four}
\hat{f}(\xi) = F[f](\xi) := \int_{\mathbb R^d} e^{ 2 \pi i \ (\xi,x)
} \ f(x) \  dx,
\end{equation}
where $  \  (\xi,x) \ $ denotes the inner (scalar) product and $ \  |x| := \sqrt{(x,x)}.  \ $ \par

Since  $f \in L_1(\mathbb R^d)$, the integral in \eqref{Four}
converges uniformly and it represents a uniformly continuous bounded
function on the whole space $\mathbb R^d$

$$
\sup_{\xi \in \mathbb R^d}|F[f](\xi)| \le |f|_1 \le ||f||G\psi_1 <
\infty.
$$

Here $\psi_1 = \psi_1(p), \ p \in [1,\infty)$ is the function
defined by \eqref{gaussian_psi}, with $\sigma=1$. For this purpose
one can replace $ \ \psi_1(\cdot) \ $ with many other such
functions, for instance, with the more general $ \
\psi_{\sigma}(\cdot), \ \sigma \in (0,\infty). \ $
\par

Recall that if $h(x) = (f * g)(x) $, then

\begin{equation} \label{FourConv}
F[h](\xi) = F[f](\xi) \cdot F[g](\xi).
\end{equation}

Let $\eta$ be an arbitrary fixed point of the  space $\mathbb R^d$;
define the set

\begin{equation} \label{IdealOrd}
J(\eta) = \{ f \,:\, f \in G\psi_1, \ F[f](\eta) = 0  \}.
\end{equation}

It follows immediately from the  relation \eqref{FourConv} that the
set $J(\eta)$ forms a well-known classical (maximal) ideal in the
Banach algebra $G\psi_1$. \par

\vspace{4mm}

Denote by $ \ UC = UC(\mathbb R^d) \ $ the set of all {\it
uniformly} continuous bounded functions $ \ f: \mathbb R^d \to
\mathbb R \ $ or $ \ f: \mathbb R^d \to \mathbb C. \ $ This set
forms also an ideal in this algebra. Namely, define the usual
modulus of continuity for a function $ \ f \in UC(\mathbb R^d) \ $

$$
\omega[f](\delta) \stackrel{def}{=} \sup_{|x-y|\le \delta} |f(x) -
f(y)|, \  \ \delta \in [0,\infty)
$$
so that
$$
f \in  UC(\mathbb R^d) \ \Rightarrow \ \lim_{\delta \to 0+}
\omega[f](\delta) = 0.
$$
For $ f \in  UC(\mathbb R^d)$ the function $x \to h(x) = (g*f)(x)$
is bounded and uniformly continuous, in fact if
$$
x_1,x_2 \in \mathbb R^d \ : \   |x_1 - x_2|\le \delta
$$
we have
\begin{equation*}
\begin{split}
|h(x_1) - h(x_2)| \le & \int_{\mathbb R^d} |g(y)| |f(x_1 - y) -
f(x_2 - y)| \, dy \\
\le & |g|_1 \omega[f](\delta) \le ||g||_{G\psi_1}  \cdot
\omega[f](\delta) \to 0, \ \ \delta \to 0+.
\end{split}
\end{equation*}
Thus, $h(\cdot) \in UC(\mathbb R^d)$.

\vspace{5mm}

{\bf D.} Let us show  now that the condition \eqref{Positive
norming} is, in the general case, essential. Namely, suppose
temporarily that $ \ \psi(1) = 0, \ $ then

$$
 \forall f \in G\psi \ \Rightarrow \ |f|_1 = 0
$$
and, consequently, $ f = 0$, i.e.  $G\psi = \{ \ 0 \ \}$.
\par
Furthermore, let us consider the following counterexample. Choose $
G = \mathbb R$ with the usual convolution. Define the function

$$
f(x)=f_{3/2}(x):= x^{-2/3} \, I(x \ge 1),
$$
where $I(A)$ denotes the indicator function of the predicate $A$.
\par

We deduce, after simple calculations,
\begin{equation*}
\tilde{\psi}(p) := |f|_p =\left\{
  \begin{array}{ll}
    \left( \displaystyle \frac{3}{2p-3}
\right)^{1/p}, & p > 3/2 \\ \\
    +\infty, & p \in [1,3/2].
  \end{array}
\right.
\end{equation*}
Evidently,
$$
||f||_{G\tilde{\psi}} = 1,
$$
as long as the function $\tilde{\psi}(p)$ is the natural function
for $f$. \par

Let us now evaluate the convolution
$$h(x) = h(x)=(f*f)(x)$$
as $ \ x \to \infty$; of course, $h(x) = 0, \ x < 1$.
 For $x \geq 3$ we have, as $ x \to \infty$,

\begin{equation*}
\begin{split}
h(x) = & \int_1^{x-1}  y^{-2/3} \ (x-y)^{-2/3} \ dy = x^{-1/3}
\int_{1/x} ^{1 - 1/x} \ z^{-2/3} \ (1 - z)^{-2/3} \  dz  \\
\sim & B(1/3, 1/3) \ x^{-1/3} = C_1 \ x^{-1/3},
\end{split}
\end{equation*}
where $B(\cdot, \cdot)$ denotes the ordinary Beta-function and $ \
C_1 = B(1/3, 1/3) \in (0,\infty)$.  \par

More precisely note that, as $x \ge 3$,

\begin{equation*}
h(x) \ge x^{-1/3} \int_{1/3}^{2/3} z^{-2/3} \ (1 - z)^{-2/3} \ dz  =
C_2 \ x^{-1/3}, \ \  C_2 = \const \in (0,\infty);
\end{equation*}

$$
|h|_p^p \ge C_2^p \int_{1/3}^{2/3} x^{-p/3} \ dx =  C_2^p
\frac{3}{p-3} \cdot \frac{1 - 2^{1 - p/3}}{3^{1 - p/3}} =:
 K^p(p), \ \ \ K(p)\in (0,+\infty);
$$

$$
|h|_p \ge K(p), \  \ p > 3,
$$

$$
 |h|_p = \infty,  \ 1 \le p \le 3.
$$

 It remains to note that the function $ \ h(\cdot) \ $ does not belongs to the Grand Lebesgue Space  $ G\tilde{\psi}$. In fact,

$$
||h||_{G\tilde{\psi}} = \sup_{p > 3/2} \left\{
\frac{|h|_p}{\tilde{\psi}(p)} \ \right\}   \ge
\frac{|h|_2}{\tilde{\psi}(2)} = \infty,
$$
or equivalently
 $$
\infty = ||f*f||_{G\tilde{\psi}} > ||f||_{G\tilde{\psi}} \cdot
||f||_{G\tilde{\psi}} = 1.
$$

Thus, the (Banach) Grand Lebesgue Space $ \ G\tilde{\psi} \ $ is not
a Banach algebra. \par

The case when
$$
f(x) = f_{\alpha}(x) = x^{-1/\alpha} \ I(x \ge 1), \ \ \alpha =
\const \in (1,2)
$$
may be investigated quite analogously. \par

By the way note that, despite $f_{\alpha}(\cdot) \notin L_1(R)$, its
Fourier transform $ F[f_{\alpha}](\xi)$ there exists as an element
of the space $L_2(R)$ because $f_{\alpha}(\cdot) \in L_2(R)$, by the
Theorem of  Plancherel-Parseval. Furthermore,
 $ F[f_{\alpha}](\xi) $ there exists for all non-zero values of the variable $ \xi$. More exactly, as $ \xi \to 0+ $

$$
F[f_{\alpha}](\xi) \sim \xi^{1/\alpha - 1}\cdot \int_0^\infty
e^{2\pi i y} y^{-1/\alpha}\,dy
$$
and

$$
\xi \to \infty \ \Rightarrow \ F[f_{\alpha}](\xi) \sim \frac{e^{2
\pi i \xi}}{2 \pi i \xi}.
$$

\vspace{6mm}

 {\it An open and interesting problem, by our opinion, is to find the necessary conditions imposed on the generating function
 $ \ \psi(\cdot) \ $  under which $ \ G\psi  = G\psi_G \ $  forms a Banach algebra and to find, in this
case, all its  maximal ideals.} \par

\vspace{6mm}

\emph{Acknowledgement.} {\footnotesize The first author has been
partially supported by the Gruppo Nazionale per l'Analisi
Matematica, la Probabilit\`a e le loro Applicazioni (GNAMPA) of the
Istituto Nazionale di Alta Matematica (INdAM) and by Universit\`a
degli Studi di Napoli Parthenope through the project \lq\lq sostegno
alla Ricerca individuale\rq\rq .\par

\end{document}